1.    A sample of material weighing m = 3.400 g was incinerated inside a bomb calorimeter. The calorimeter was charged with mW = 1500 g of water. The initial temperature of the calorimeter was t0 = 24.223°C. After the incineration the temperature rose to t1 = 29.654°C.

Calculate the heat of combustion of 1 kg of tested material.

Heat capacity of the calorimeter: cP,(c) = 3.3 kJ/K

Specific heat of water: CP,(W) = 4.18 kJ/(kg·K)

1
Expert's answer
2022-02-17T08:16:03-0500

Q= m•c•dT

dT= t1-t0= 29.654-24.223= 5.431°c+ 273.15

dT= 278.581K

Q= 1× 3.3× 278.581= 919.32kJ